OpenStudy (anonymous):

X/8+6=-12 ; -48 tell me if its a solution and show the steps thanks?

jimthompson5910 (jim_thompson5910):

x/8+6=-12 (-48)/8+6=-12 ... replace x with -48 -6+6=-12 0=-12 Since the equation is NOT true, this means that -48 is NOT a solution.

OpenStudy (anonymous):

\[\frac{-48}{8} + 6 \neq 12\] There fore it is not a solution

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

Therefore***

OpenStudy (anonymous):

how did you get -6+6=-12

OpenStudy (anonymous):

nvm

jimthompson5910 (jim_thompson5910):

-48/8 = -6

OpenStudy (anonymous):

 pizza!(: